29.) 6 points. Consider the following data concerming the equation H:02+31 +2H +210 rate H202] [H1 1 0.100 M 5.00x 10 M 1.00 x 102 M 0.137 M/sec II. 0.100 M 1.00 x 10-3 M 1.00 x 102 M 0.268 M/sec 111. 0.200 M 1.00x10-) M 1.00x10-2 M 0.542 Msec IV. 0.400 M 1.00 x 10-3 M 2.00 ×10-2 M 1.084 Msec What is the rate law and the value of k for this reaction? a. ,-) //0.100AI K?S. 00x {? s, ook Two mechanisms are proposed: Explain which mechanism and which step as the rate determining step would best fit the data? b. meure h+11 Stee
Answer #1

a). To write the rate law expression first of all we have to calculate the order of the reaction. For this we will write the rate expression by using the given data and then we will calculate the order with respect to each reactant.

Rate = k[H2O2]^a × [I-]^b ×[H+]^c

Here a,b,c are order for the reaction with respect to reactants accordingly.

FROM EXPERIMENT 1.

0.137 M/s = k(0.100M)^a × (5.0×10^-4M)^b × (1.00×10^-2M)^c

FROM EXPERIMENT 2.

0.268M/s = k(0.100M)^a × (1.0×10^-3M)^b × (1.00×10^-2M)^c

FROM EXPERIMENT 3.

0.542M/s = k(0.200M)^a × (1.0×10^-3M)^b ×(1.00×10^-2M)^c

FROM EXPERIMENT 4.

1.084M/s = k(0.400)^a×(1.0×10^-3)^b × (2.0×10^-2)^c

espert 68)